### v1

# On the Structure and Semantics of Identifier Names Containing Closed Syntactic Category Words

Abstract
Identifier names are crucial components of code, serving as primary clues for developers to understand program behavior. This paper investigates the linguistic structure of identifier names by extending the concept of grammar patterns; representations of the part-of-speech (PoS) sequences that underlie identifier phrases. The specific focus is on closed syntactic categories (e.g., prepositions, conjunctions, determiners), which are rarely studied in software engineering despite their central role in general natural language. The Closed Category Identifier Dataset (CCID) is presented, a new manually annotated dataset of 1,275 identifiers drawn from 30 open-source systems. The relationship between closed-category grammar patterns and program behavior is analyzed using grounded theory coding, statistical, and pattern analysis. The results reveal recurring structures that developers use to express control flow, data transformation, temporal reasoning, and behavioral roles through naming. This study contributes an empirical foundation for understanding how developers adapt linguistic resources to encode behavior in source code. By analyzing closed-category terms and their associated grammar patterns, the work highlights a previously underexplored dimension of identifier semantics and identifies promising directions for future research in naming support, comprehension, and education.
Keywords:
identifier naming program comprehension part of speech tagging software maintenance and evolution software linguistics closed category terms naming conventions
journal: Empirical Software Engineering
1 Introduction
Developers spend a significant amount of time reading and comprehending code Corbi1989 ; Martin:2008 , and identifier names play a central role in this process, accounting for roughly 70% of all code characters Deissenbock:2005 . Prior work shows that the quality of identifier names significantly impacts comprehension Schankin:2018 ; Binkley2006 ; Hofmeister:2017 ; butler2010exploring ; Takang1996 ; feitelson:2017 ; fakhoury:2020 , supports tooling Binkley:2018 ; newmanabbrev , and poses persistent pedagogical challenges Felienne:2024 ; Glassman2015FoobazVN . These challenges motivate research into how naming practices encode meaning, and how we might better characterize or improve them.
A key obstacle in studying identifier names is measuring the semantics they convey, not just at the level of individual terms, but in the structure and composition of entire names. Some approaches cluster identifiers by terms or embeddings Allamanis:2015 ; Liu:2019 , while others analyze them using syntactic or static roles Dragan:2006 ; Alsuhaibani:2015 ; Newman:2017 . In this work, we focus instead on grammar patternsNewman_GP : sequences of part-of-speech (PoS) tags that abstract the phrasal structure of identifiers. Grammar patterns provide a syntactic lens through which naming semantics can be studied at scale, offering insight into how term combinations convey behavioral meaning.
At a high level, PoS can be split into two Syntactic Categories: open and closed. Most identifier naming research has focused on open-category, which includes nouns and verbs. The set of open category terms changes and increases (in terms of new words) with time as new domains emerge and change. In contrast, closed-category (e.g., prepositions, conjunctions, determiners) are drawn from a fixed set and serve functional roles in language; this set of terms rarely sees new words introduced over time. These terms have received little attention in the software literature, despite their importance in human languages. Identifying closed-category terms in code is also non-trivial: for example, the word and may represent a conjunction or a logical operator, depending on context—making PoS tagging a prerequisite for meaningful analysis.
The goal of this paper is to investigate how closed-category terms are used in identifier names to express program behavior, using the grammar patterns (see Section 3 for definitions) that these terms appear within to provide insights into how these terms interact with the other terms around them. We extend prior research on general grammar patterns Newman_GP ; newman2021Catalogue by introducing and analyzing the Closed Category Identifier Dataset (CCID), a manually annotated corpus of 1,275 identifiers from 30 open-source systems. Unlike raw term-based approaches, grammar patterns abstract away surface vocabulary, allowing us to characterize naming conventions by their syntactic structure. By examining both the patterns and the concrete terms that instantiate them, we explore how developers use compact linguistic forms to encode behavioral semantics in code. Specifically, we contribute:
  • A new dataset (CCID) of identifiers containing closed-category terms, annotated with PoS tags, grammar patterns, and contextual metadata.
  • A mixed-methods analysis combining grounded theory coding with statistical evaluation to characterize the semantics of closed-category grammar patterns and their constituent terms.
  • An evaluation of how these patterns correlate with programming context, language, and domain.
Our findings have implications for both human and automated naming support. Grammar patterns offer a structured way to analyze naming behavior, surface potential inconsistencies, and guide naming suggestions. For AI-based tools, they offer scaffolding to align generated names with human conventions. For developers and educators, they reveal naming idioms that can support clearer communication and pedagogy. In this study, we address the following research questions:
RQ1: What behavioral roles do closed-category terms play in source code identifiers? To address this question, we conduct a grounded theory study on a manually annotated dataset of identifiers containing closed-category terms: prepositions, conjunctions, determiners, and digits. Through iterative coding and memoing, we develop axial and selective codes that describe the behavioral functions these terms convey in source code, such as data flow, condition handling, or execution sequencing. This process allows us to uncover not only common grammar patterns but also the communicative intent behind developers’ use of closed-category terms. Our goal is to characterize the nuanced and purposeful ways in which these terms encode program behavior.
RQ2: How do closed-category terms correlate with structural, programming language, and domain-specific contexts in software? To answer this question, we quantitatively analyze the distribution of closed-category terms across multiple dimensions: source-code-local structure (e.g., function names, parameters, class names), programming languages (e.g., Java, C++, C), and system domains (e.g., libraries, frameworks, domain-specific applications). We use statistical tests to examine whether these terms appear disproportionately in certain contexts. These correlations help us determine whether developers systematically leverage closed-category terms to express behavior in ways that are shaped by structural conventions, linguistic norms, or domain constraints.
This paper is organized as follows. Section 2 provides our reasoning on why it is important to study this topic. Section 3 gives background on grammar patterns in the context of identifier names. Section 4 provides a detailed explanation of our methods for undertaking the investigation. Our Evaluations are presented in Sections 5 and 6. Related work on identifier names is in Section 7. Discussion of the results is in section 8, followed by Threats to Validity in Section 9. Conclusions are in Section 10 and Data Availability in Section 11.
2 Why Study Closed-Category Naming Patterns?
Closed-category terms are relatively uncommon in identifier names. Because they are uncommon, their presence raises an important question: When developers do use these terms, what specific meaning or behavior are they trying to convey? We hypothesize that developers include closed-category terms deliberately, as a way to encode behaviorally specific semantics that are lost or obscured without them. Consider the following examples:
  • find_all_textures: The determiner all signals a universal scope, clarifying that this identifier refers to the entire set of textures, not a subset.
  • on_start: The preposition on reflects event-driven logic, indicating that the associated behavior is triggered at the start of execution.
  • warn_if_error: The conjunction if embeds a conditional relationship, revealing that the action is contingent on an error occurring.
In each case, the closed-category term is essential to understanding the behavioral semantics of the identifier. Without these terms, the names are more ambiguous or less informative. While uncommon in aggregate, closed-category terms often signal precise intent and encode logical structure in compact forms.
Despite their potential significance, these terms have received almost no attention in prior software development naming research, which has focused primarily on open-category words (e.g., nouns, verbs). As a result, we lack foundational knowledge about when and how closed-category terms are used in code—and what they contribute to program comprehension.
Understanding these naming patterns has clear implications: it can inform naming tools, guide educational resources, improve automated name generation, and help researchers characterize naming conventions more precisely. Closed-category terms may be uncommon, but we argue, in this paper, that their usage is not accidental; they significantly contribute to the meaning of identifier names, making it important to study them.
Table 1: Examples of closed-category grammar patterns
Identifier Example Grammar Pattern
action to index map N P NM N
as binary P N
time for each line N P DT N
server and port N CJ N
open if empty V CJ NM
adjust to camera V P N
Table 2: Part-of-speech categories used in study
Abbreviation Expanded Form Examples
N noun
stack, function, language
DT determiner
the, this, that, these, those, which
CJ conjunction and, for, nor, but, or, yet, so
P preposition
behind, in front of, at, under,
beside, above, beneath, despite
NPL noun plural
strings, identifiers, classes
NM (bold) noun modifier
employeeName, tokenParser
V verb run, execute, implement, develop
VM verb modifier (adverb)
quickly, safely, eventually
PR pronoun
she, he, her, him, it, we,
us, they, them, I, me, you
D digit 1, 2, 10, 4.12, 0xAF
PRE preamble* Gimp, GLEW, GL, G
Figure 1: Examples of noun, verb, and prepositional phrases
3 Definitions & Grammar Pattern Generation
In this work, we analyze identifier names through the lens of grammar patterns, which are sequences of part-of-speech (PoS) tags assigned to the terms within an identifier. For example, the identifier GetUserToken is split into the terms Get, User, and Token, which are tagged as Verb Noun-adjunct Noun. This sequence, V NM N, represents the identifier’s grammar pattern. Crucially, this pattern generalizes across many identifiers: RunUserQuery and WriteAccessToken share the same structure, even though they use different terms. Grammar patterns thus allow us to relate identifiers by their syntactic form.
We focus specifically on closed-category grammar patterns, which are patterns that contain at least one closed-class part of speech: a preposition, determiner, conjunction, or numeral (digit). These categories are finite and rarely accept new terms, in contrast to open-class categories like nouns and verbs, which grow over time as new domains introduce new concepts. Despite their rarity in code, closed-category terms often signal behavioral relationships such as event triggers, quantification, or conditional logic, making them important to study.
Part-of-Speech Tags. Table 2 lists the PoS tags used in this study. Most are drawn from standard linguistic categories. We highlight one custom tag that is central to our analysis:
  • Noun Modifier (NM): Includes adjectives as well as noun-adjuncts—nouns used to modify another noun (e.g., user in userToken, or content in contentBorder). Although standard PoS taggers do not typically distinguish noun-adjuncts, prior work shows their critical role in naming semantics Newman_GP .
  • Preamble (PRE): A prefix used to convey structural or language-specific metadata, rather than domain semantics. Common examples include Hungarian-style markers such as m_ for member variables, or project-level namespaces like gimp in gimp_temp_file; a practice especially common in C. For a complete typology and discussion, see Newman_GP ; we include preambles here since we do use them in the data set, but they are not the focus of this paper.
3.1 Phrasal Structures and Interpretation
While our analysis is based on PoS sequences rather than full parse trees, we draw on linguistic phrase structure to interpret identifier patterns. Specifically, we reference three example concepts to help the reader understand what we mean when we use the term ‘phrase’ with respect to grammar patterns:
  • Noun Phrase (NP): A noun optionally preceded by one or more modifiers (e.g., accessLog, userToken, windowTitle).
  • Verb Phrase (VP): A verb followed by a noun phrase, often representing an action on a specific entity (e.g., getUserToken, drawContentBorder).
  • Prepositional Phrase (PP): A preposition followed by a noun phrase (e.g., onClick, fromCache).
These phrase structures help illustrate how grammar patterns support analysis of phrases. For instance, in drawContentBorder, the noun-modifier content refines the meaning of the head noun border, while the verb draw anchors the identifier as a behavior applied to that concept (i.e., draw applied to a specific type of border; a content-border). When closed-category terms appear, they may indicate when an action should occur (onStart), under what condition (ifError), or which entities are included (allTextures). Figure 1 shows examples of NP, VP, and VP-with-PP constructions as derived from grammar patterns.
Table 3: List of 30 open source systems included in study
Repo Link Name Primary Language Commit Hash Data of most recent commit C LOC C++ LOC Java LOC Total LOC
https://github.com/liuliu/ccv ccv C 3eb1664f88e04515bf6861663c4f96d6279a47f5 2023-07-19 279186 1908 0 281094
https://github.com/ropensci/git2r git2r C 9c42d41c14cf83735db37ae662e2bce42c500aaa 2023-05-01 99956 0 0 99956
https://github.com/Juniper/libxo libxo C 290d88625c683bca60e05e68dd2003c290f50f99 2023-02-08 9361 0 0 9361
https://github.com/mgba-emu/mgba mgba C c8cfaefcc880e668b0db8ba5f8b49978f0d108fd 2023-07-18 302865 25309 0 328174
https://github.com/naemon/naemon-core naemon-core C 7558c98937ffb15c040daf158f92827b1f338ba6 2023-07-07 40991 0 0 40991
https://github.com/openvswitch/ovs ovs C f5188ff2147517612b410ed607e3843cdf4b51a6 2023-07-19 268376 0 0 268376
https://github.com/igraph/rigraph rigraph C cae1f110975e38353a8fbbe16d4c7a008fbd9bea 2023-07-19 273973 32737 0 306710
https://github.com/toggl-open-source/toggldesktop toggldesktop C 013038304ee8e9154db4a45fb642716956cf153b 2023-06-22 582087 269105 0 851192
https://github.com/irungentoo/toxcore toxcore C bf69b54f64003d160d759068f4816b2d9b2e1e21 2018-10-03 27443 0 0 27443
https://github.com/weechat/weechat weechat C 13241cdca5b39415d2a0f9e1b55a79153d14ba07 2023-07-20 197608 31175 0 228783
https://github.com/wireshark/wireshark wireshark C 13f21ebd59317f59b91ce7a48a234ef48989ccf0 2023-07-20 4171790 102848 0 4274638
https://github.com/BVLC/caffe caffe C++ 9b891540183ddc834a02b2bd81b31afae71b2153 2020-02-13 0 42856 0 42856
https://github.com/vgvassilev/cling cling C++ 5e3310ac045d5763cd9852359d1003f3850c2a0c 2023-07-18 57 28342 0 28399
https://github.com/ipkn/crow crow C++ 2b43d3cd6a9a9cdbc99dfef9b86ff3f3027f3d1f 2022-09-20 0 1434 0 1434
https://github.com/fakeNetflix/facebook-repo-ds2 ds2 C++ eaf83c59b8fff11cc4c218da7b78a983754727ec 2019-07-17 367 27011 0 27378
https://github.com/freeminer/freeminer freeminer C++ 86caf1cfbdf3d06f53981adaedc6f1eba7448486 2023-04-22 15090 130828 1077 146995
https://github.com/meta-toolkit/meta meta C++ c7019401185cdfa15e1193aad821894c35a83e3f 2017-08-19 132 25451 0 25583
https://github.com/panda3d/panda3d panda3d C++ 07f9f9d89722a82234e0d111e61b7ff98e88e2db 2023-06-13 44671 416212 175 461058
https://github.com/facebook/proxygen proxygen C++ 5370e58851d6b363f1edbc542a0daa369ebbf303 2023-07-20 2676 70161 0 72837
https://github.com/s3fs-fuse/s3fs-fuse s3fs-fuse C++ 5371cd1468c84423729c334ac997f9621b797e9f 2023-07-19 197 19582 0 19779
https://github.com/cglib/cglib cglib Java 5c947539c1218dc4ff1a14a7c36b4856b0ee6397 2022-02-08 0 0 15187 15187
https://github.com/deeplearning4j/deeplearning4j deeplearning4j Java 854c3b264d6d392ecd1b09c386d919e3c74cb425 2023-06-21 0 224997 696717 921714
https://github.com/apache/drill drill Java f558815feb6560ebbc3d1fcf502acee8fea1e555 2023-06-21 538 34591 626295 661424
https://github.com/google/guava guava Java 551f92130cc0442870569b89820e0b432a8ff8f7 2023-07-18 0 0 356651 356651
https://github.com/immutables/immutables immutables Java 28ba97e9f3075902e279c1cde3e16be01e365afb 2023-06-16 0 0 69505 69505
https://github.com/dropwizard/metrics metrics Java d6cf6a4e8f6e65a0b6107be6d8d73e00f0d95efa 2023-07-20 0 0 31317 31317
https://github.com/igniterealtime/Openfire Openfire Java 25ed0d66f9c56f423f9408920ce7d8734dcaf84d 2023-07-20 120 0 122186 122306
https://github.com/HubSpot/Singularity Singularity Java 15688f486fc9286878eff40b51789c88bd6899d5 2022-11-18 0 0 122183 122183
https://github.com/igniterealtime/Smack Smack Java 19b20fefeca69cccc3637e01ec2f151a054a3351 2023-04-26 0 0 125547 125547
https://github.com/igniterealtime/Spark Spark Java ca028f3a818b94c00f2f600933109087d84fb0b9 2023-05-11 9 0 91886 91895
TOTAL 6317493 1484547 2258726 10060766
4 Methodology
For our study, identifiers are collected from and analyzed in the following contexts: class names, function names, parameter names, attribute names (i.e., data members), and declaration-statement names. A declaration-statement name is a name belonging to a local (to a function) or global variable. We use this terminology because it is consistent with srcML’s terminology collard:2016 for these variables, and we used srcML to collect identifiers. Therefore, to study closed-category grammar patterns, we group identifiers based on these five categories. The purpose of doing this is to study closed-category grammar pattern frequencies based on their high-level semantic role (e.g., class names have a different role than function names). We collected these identifiers from 30 open source systems, which can be found in Table 3. These systems belonged to a curated dataset of engineered software projects, synthesized by Reaper Munaiah2017 , which is a tool that measures how well different projects follow software engineering practices such as documentation and continuous integration.
The set of systems have an average and median of 335,358 and 111,069 LOC, respectively. 11 of the systems are primarily C systems, 9 are primarily C++, and 10 are primarily Java. We chose systems that have tests and use continuous integration (CI) under the idea that these represented systems that have at least some basic process for ensuring quality; Reaper is able to automatically determine which systems have both CI and tests. Our primary concern for selecting systems is that they represent different programming languages, follow basic quality procedures, and are large enough for us to collect enough identifiers. Given this, our choice of systems is designed to ensure that the grammar patterns in this study are applied across at least the languages under study.
Table 4: Distribution of part-of-speech labels in Old Data Set and CCID
Old Data Set CCID
TAG FREQUENCY TAG FREQUENCY
NM 1604 (45.2%) N 1141 (31.58%)
N 1141 (32.1%) NM 643 (17.79%)
V 305 (8.6%) P 398 (11.02%)
NPL 238 (6.7%) V 363 (10.04%)
PRE 105 (3%) DT 308 (8.52%)
P 94 (2.6%) D 283 (7.83%)
D 27 (0.8%) PRE 217 (6.00%)
DT 15 (0.4%) NPL 142 (3.93%)
VM 13 (0.4%) VM 69 (1.91%)
CJ 8 (0.2%) CJ 50 (1.38%)
Total 3550 Total 3614
Table 5: Distribution of Tags in Candidate and Verified (Manually-annotated) data set
CJ DT
Candidate Verified Candidate Verified
Attribute 66 (28.09%) 6 (12.24%) 78 (24.92%) 84 (27.54%)
Declaration 62 (26.38%) 10 (20.41%) 79 (25.24%) 85 (27.87%)
Parameter 44 (18.72%) 6 (12.24%) 78 (24.92%) 58 (19.02%)
Function 63 (26.81%) 27 (55.10%) 78 (24.92%) 78 (25.57%)
Class 0 (0.00%) 0 (0.00%) 0 (0.00%) 0 (0.00%)
Total 235 49 313 305
D P
Candidate Verified Candidate Verified
Attribute 80 (21.98%) 62 (23.40%) 89 (24.45%) 103 (26.96%)
Declaration 81 (22.25%) 70 (26.42%) 88 (24.18%) 73 (19.11%)
Parameter 81 (22.25%) 77 (29.06%) 89 (24.45%) 60 (15.71%)
Function 80 (21.98%) 41 (15.47%) 88 (24.18%) 140 (36.65%)
Class 42 (11.54%) 15 (5.66%) 10 (2.75%) 6 (1.57%)
Total 364 265 364 382
Table 6: Balanced population of identifiers per context
Context Sample Population
Attribute 312 (24.47%)
Declaration 306 (24.00%)
Function 313 (24.55%)
Class 52 (4.08%)
Parameter 292 (22.90%)
Total 1275
4.1 Detecting and Sampling Identifiers with Closed-Category Terms
Sampling identifiers that contain closed-category terms is challenging for two reasons: (1) They are relatively uncommon in production code, and (2) many such terms are ambiguous without context, making automatic tagging difficult. To address this, we implemented a two-phase sampling strategy: (1) filtering identifiers that potentially contain closed-category terms into candidate sets, and (2) manually verifying and annotating a statistically representative sample.
Phase 1: Filtering Candidate Identifiers
We began with the CCID corpus, which contains 279,000 unique identifiers from production code. To collect these 279K identifiers, we used the srcML identifier getter tool  [^1] https://github.com/SCANL/srcml_identifier_getter_tool on the srcML archives resulting from running srcMLcollard:2016 on the system repository directories (Table 3). To identify candidate sets:
  • Digits (D): We selected identifiers containing at least one numeric digit, using Python’s isDigit() functionality. Digits are easier to detect automatically and unambiguous in token form. However, there are cases where a digit will be annotated as part of another category. For example, str2int uses the digit 2 as a preposition (to).
  • Determiners (DT), Conjunctions (CJ), and Prepositions (P): We constructed lexicons for each category using curated lists of common English terms [^2] https://en.wikipedia.org/wiki/List_of_English_determiners [^3] https://www.englishclub.com/grammar/prepositions-list.php [^4] https://7esl.com/conjunctions-list/. We then filtered for identifiers containing component words (i.e., split tokens) that matched a word in one of these lists. This approach is viable only because these categories are closed and finite in vocabulary.
This filtering process produced the following candidate counts:
  • 602 identifiers candidate conjunctions
  • 1,693 identifiers candidate determiners
  • 3,383 identifiers candidate prepositions
  • 4,630 identifiers candidate digits
We use the term candidate because these filters do not account for context or usage, and thus include false positives. Still, they serve as an upper bound for each category’s prevalence in the corpus. Based on this, we estimate the proportion of identifiers containing each term type as follows:
  • 0.2% (602/279,000) contain conjunctions
  • 0.6% (1,693/279,000) contain determiners
  • 1.2% (3,383/279,000) contain prepositions
  • 1.7% (4,630/279,000) contain digits
Phase 2: Balanced Sampling and Annotation
Using a 95% confidence level and a 5% margin of error, we computed minimum sample sizes for each category. For example, a 95 and 5 sample for conjunctions (602 identifiers) is 235:
  • CJ: 235 candidate conjunction identifiers
  • DT: 313 candidate determiner identifiers
  • P: 345 candidate preposition identifiers
  • D: 355 candidate digit identifiers
Before manual annotation, we stratified the candidate identifiers by their program context:
  • Function names
  • Parameters
  • Attributes (i.e., class members)
  • Function-local declarations
  • Class names
Some contexts, such as parameters and especially class names, were underrepresented due to the natural scarcity of closed-category terms in those positions. To increase representation from underrepresented contexts, we attempted to oversample relevant subgroups. However, even with oversampling, the absolute number of qualifying identifiers (e.g., a parameter containing a conjunction) remained low. As our sampling was driven by the presence/population of closed-category terms in general, rather than population within our program contexts under study, we opted not to artificially balance the dataset further.
After sampling and stratification, we obtained a total of 1,275 identifiers across the four categories in our candidate set:
  • 364 candidate preposition identifiers
  • 363 candidate digit identifiers
  • 313 candidate determiner identifiers
  • 235 candidate conjunction identifiers
Following manual verification and annotation (described in Section 4.2), we retained only those identifiers that were confirmed to contain closed-category terms. Table 5 summarizes both the sampled totals and the verified counts. The final dataset consists of 1,001 identifiers confirmed to contain at least one closed-category term. These identifiers comprise the CCID. Table 6 shows the CCID, but broken down by program context instead of closed-category type.
Annotation Notes
Some tools exist for part-of-speech tagging of source code identifiers (e.g., the ensemble tagger from newman_ensemble ), but these are slow at scale and are trained on datasets that underrepresent closed-category terms. For example, in our prior dataset Newman_GP , used to train the aforementioned tagging approach newman_ensemble ; conjunctions, determiners, and prepositions made up only 0.2%, 0.4%, and 2.6% of tags, respectively. Thus, we found manual annotation necessary to ensure sufficient coverage and correctness for our study.
As we are primarily concerned with production code, and prior work shows that test name grammar patterns differ from production names peruma_test_name_gp , we did not collect any identifier containing the word ‘test’, or that appeared in a clearly marked test file or directory. In addition, note that Table 4 counts tags at the word level (e.g., CJ CJ N counts two CJ tags), whereas Table 5 counts tags at the identifier level (e.g., one identifier with multiple CJ tags counts as one). This explains occasional mismatches between sampled and actual tag distributions.
4.2 Manual Process for Annotating Part-of-Speech
Initially, one author (annotator) is assigned annotate each identifier in the CCID with its grammar pattern. The annotator has experience annotating identifiers with PoS from prior work Newman_GP ; peruma_test_name_gp . The process is as follows: The annotator is given a split (using Spiral HuckaSpiral ) identifier along with the identifier’s type, file path, and line number to make it easy to find the identifier in the original code. The annotator is allowed to look at the source code from which the identifier originated if necessary. The annotator is asked to additionally identify and correct mistakes made by Spiral. When the annotator is finished, two additional annotators are asked to validate (agree or disagree) with the annotations created by the original annotator. Any disagreements are discussed and fixed, if required. Furthermore, a fourth annotator made annotations, which are then compared to the original annotator’s work. Again, disagreements are discussed and fixed. An example disagreement is with the identifier where_len, which is a tricky one because ‘where’ is typically an adverb or conjunction. However, in this case ‘where’ is a reference to a void pointer variable called ‘where’ within the code. So ‘where_len’ is the length of the memory this pointer points to, making ‘where’ a noun-adjunct in this case; describing the type of length. Thus, its grammar pattern is NM N. The Fleiss’ Kappa for this process was .916.
We did not expand abbreviations for a couple reasons. The first is that some abbreviations are more meaningful than their expanded terms (e.g., HTTP, IPV4, SSL) due to how frequently they are used in their abbreviated form by the community. The second reason is that abbreviation expansion techniques are not widely available and vary widely in terms of effectiveness on different types of terms newmanabbrev ; Jingxuan:2023 . Therefore, a realistic worst-case scenario for developers and researchers is that no abbreviation-expansion technique is available to use, and their PoS taggers must work in this worst-case scenario. Whenever we recognized one, we do not split domain-term abbreviations (e.g., Spiral will make IPV4 into IPV 4; we corrected this to IPV4). We do this because it is the view of the authors that they should be recognized and appropriately tagged in their abbreviated (i.e., their most common) form.
Table 7: Most common Patterns
Determiner Conjunction Preposition Digit
DT N 109 (35.74%) N CJ N 6 (12.24%) P N 70 (18.32%) N D 125 (47.17%)
DT NM N 40 (13.11%) V CJ N 3 (6.12%) P NM N 31 (8.12%) NM N D 33 (12.45%)
DT NPL 20 (6.56%) CJ NM 2 (4.08%) P V 13 (3.40%) N D N 9 (3.40%)
DT V 7 (2.30%) NM CJ NM 2 (4.08%) N P N 12 (3.14%) PRE N D 8 (3.02%)
DT NM NM N 6 (1.97%) V CJ V 2 (4.08%) V P N 12 (3.14%) V N D 8 (3.02%)
N DT 6 (1.97%) V N CJ N 2 (4.08%) P 10 (2.62%) NPL D 6 (2.26%)
V DT 6 (1.97%) CJ 1 (2.04%) NM N P N 9 (2.36%) N D NM N 3 (1.13%)
DT NM NPL 5 (1.64%) CJ DT N 1 (2.04%) V P 9 (2.36%) V NM N D 3 (1.13%)
V DT N 5 (1.64%) CJ NM N 1 (2.04%) N P 8 (2.09%) N D NPL 2 (0.75%)
DT 4 (1.31%) CJ V 1 (2.04%) NPL P N 8 (2.09%) N P D NPL 2 (0.75%)
5 Evaluation of RQ1: What behavioral roles do closed-category terms play in source code identifiers?
Our evaluation aims to establish, through RQ1 and RQ2: 1) how closed-category terms are used to convey differing types of program behavior, 2) the typical grammatical structure of identifiers containing closed-category terms, and 3) how closed-category term distributions differ across programming context, language, and system domains. This research question investigates the semantic role of closed-category grammatical patterns in identifier naming. We focus on four closed-category part-of-speech types: prepositions, digits, determiners, and conjunctions. We present our findings by (1) describing each category’s semantic function using axial codes, (2) summarizing behavioral trends via selective coding, and (3) highlighting shared trends through cross-category synthesis.
5.1 Methodology: Manual Process for Grounded Theory Annotations
We employed a grounded theory approach to analyze variable names in source code and their relationship to program behavior. This multi-phase coding process involved four annotators and combined individual annotation with iterative validation and synthesis to construct a theory grounded in observed naming patterns. The sample used in this study is a subset of the CCID described in Section 4. To construct this subset, we took the top 10 most common grammar patterns (Table 7) and collected all identifiers that followed these patterns; randomly selecting the 10th grammar pattern if multiple patterns had the same frequency. These represent the most common names (i.e., from the perspective of grammatical structure) used in the data set. This totals to 618 identifiers.
Four annotators participated in the process, comprising both faculty and graduate students with prior experience in natural language processing and software engineering research. All annotators had previously worked on part-of-speech annotation tasks. Before formal annotation began, the team conducted a one-hour training and calibration session to discuss the guidelines, walk through examples, and establish expectations and deadlines.
Coding Platform.
Annotations are conducted collaboratively using a shared Google Sheets document. Each row in the sheet contained an identifier along with contextual metadata, including:
  • Identifier Name
  • Source Code Context
  • Programming Language
  • GitHub Commit Link
  • Split Identifier Name (tokenized form)
  • Grammar Pattern (POS sequence)
  • Notes (for open coding and memoing)
  • Axial Code (for grouping behavioral patterns)
Open coding and memoing are captured directly in the Notes column. The final axial codes were recorded in the corresponding column once annotators had synthesized their observations.
Phase 1: Familiarization.
All annotators reviewed the dataset to build familiarity with the variable names, associated grammar patterns, and program contexts. They discussed ambiguous or novel constructions in group chats to align interpretations and maintain consistency.
Phase 2: Open Coding.
Annotators examined each variable name in its context and assigned a free-form behavioral interpretation based on how the variable is used in the surrounding code. These open codes and rationale are documented in the Notes column. The goal was to capture a grounded understanding of what each identifier conveyed, informed by both linguistic structure and program behavior.
Phase 3: Axial Coding.
Annotators grouped similar open codes into higher-level axial codes, focusing on patterns where particular grammar structures aligned consistently with specific behavioral roles. These axial codes captured mid-level abstractions (e.g., State Variables, Event Triggers), and were documented in the spreadsheet alongside notes justifying the grouping where needed. The Fleiss’ Kappa for this phase was: .971 for Digits, .996 for Determiners, .976 for Prepositions, and 1.0 for Conjunctions.
Each annotator’s axial codes were reviewed by a different annotator for validation. This cross-review process involved reading both the open codes and the proposed axial codes, discussing disagreements, and refining the categories until consensus was reached.
Phase 4: Selective Coding.
One annotator synthesized the final, validated axial codes across all annotations and constructed a set of selective codes representing core theoretical categories that linked grammar structure to program intent. These selective codes were then shared with the remaining annotators, who were asked to evaluate whether they reflected the themes and relationships they had observed during their own coding work. Annotators agreed or suggested revisions to finalize the theory.
5.2 Digits in Identifiers
Overview.
Digits in identifiers act as compact, semantic indicators of structure, ordering, or version. They are also often used to disambiguate entities and encode numeric conventions. Their meaning is typically inferred through domain knowledge, making them powerful, but potentially hard to understand for those without the requisite domain knowledge.
Axial Codes.
We created a dual-axis framework for interpreting the meaning of digits, inspired by a single-axis framework we created in prior work on digits in identifiers peruma_digits . This framework reflects our observation that digits contribute information in two distinct ways: (1) the role they play within the local context (e.g., indexing, versioning), and (2) the source of meaning they draw from, which is often external to the immediate source code scope (e.g., domain conventions, technical standards). Every digit in the set has both a role and a source of meaning; they must be combined to fully understand the digit. We put an ‘x’ between each combination of ‘Role’ and ‘Source of Meaning’ Axial Code.
  • Role: What functional purpose the digit serves in the identifier.
    • Distinguisher: The digit differentiates conceptually similar entities, typically to avoid name collision errors from the compiler (e.g., arg1, tile2).
    • Version Identifier: The digit encodes versioning information such as protocol revisions or data format versions (e.g., http2, v1).
  • Source of Meaning: Where the interpretation of the digit originates, typically via convention, tooling, or domain-specific logic.
    • Auto-Generated: The digit is added automatically by tools, compilers, or naming systems to avoid conflicts (e.g., var1_2, jButton3).
    • Human-Named Convention: The digit’s meaning is primarily derived from ad hoc developer intent and is not more complex than distinguishing entities manually (e.g., str1, feature2).
    • Locally Specific Concept: The digit conveys project- or context-specific information, often related to coordinate systems, data structures, or memory layouts (e.g., m33 for matrix row 3 col 3).
    • Technology Term / Standard: The digit is part of a recognized domain-specific label, format, or protocol (e.g., HTTP2, Neo4j).
Role x Source of Meaning
  1. 1.
    Distinguisher × Human-Named Convention (122 items)
    Description: This group captures identifiers that use manually assigned numeric suffixes to distinguish conceptually and lexically similar entities.
    Examples: host1 (first of 2 host variables), e8 (element 8 in a parameter list)
    Grammar patterns:
    • N D (73)
    • NM N D (21)
    • V N D (6)
    • NPL D (6)
    • N D N (5)
    • PRE N D (3)
    • N D NM N (2)
    • P D (2)
    • PRE NM N D (2)
    • V NM N D (2)
  2. 2.
    Distinguisher × Locally Specific Concept (45 items)
    Description: This group captures identifiers where digits encode positional or logical roles based on system-specific conventions, such as grid layout or data structure indexing.
    Examples: dist2 (squared distance calculation), col1 (first column of a matrix)
    Grammar patterns:
    • N D (32)
    • NM N D (5)
    • PRE N D (3)
    • N D N (2)
    • NM N D P D (2)
    • V N D (1)
  3. 3.
    Distinguisher × Technology Term / Standard (17 items)
    Description: This group captures identifiers that include digits as part of standardized or domain-specific naming conventions, often encoding formats or specifications.
    Examples: b1110 (binary for UTF8 byte sequences), count32 (32-bit count value)
    Grammar patterns:
    • N D (7)
    • NM N D (5)
    • PRE N D (2)
    • N D N (1)
    • V N D (1)
    • V NM N D (1)
  4. 4.
    Version Identifier × Technology Term / Standard (9 items)
    Description: This group captures identifiers where the digit signals the version number of a protocol, tool, or technology component.
    Examples: gw6 (gateway addr for IPV6), httperf2 (version 2 of the httperf tool)
    Grammar patterns:
    • N D (5)
    • NM N D (2)
    • N D N (1)
    • N D NM N (1)
  5. 5.
    Distinguisher × Auto-Generated (8 items)
    Description: This group captures identifiers that are automatically suffixed with a digit to ensure uniqueness, often generated by tools or compilers.
    Examples: field37, field4 (numbers are generated to avoid name collissions)
    Grammar patterns:
    • N D (8)
Example.
Consider the identifier m34, which appears in the context of a matrix operation. To fully interpret the digit 3 in this name, we must consider both its Role and its Source of Meaning. Semantically, the digit serves as a Distinguisher; uniquely identifying this variable apart from its siblings (such as m32 and m31). However, its full interpretation depends on its Locally Specific Concept source: the developers have an internal convention that 3 refers to the row index, while 4 refers to the column. Without knowing the Source of Meaning, the numbers can only be interpreted as distinguishing one identifier from another; the meaning of the digits would remain ambiguous. This illustrates how both axes work together—Role tells us what the digit is doing, while Source of Meaning tells us how to interpret the value.
Selective Coding Insight.
Digits serve as semantic compression tools in source code: conveying versioning, layout, ordering, or configuration state using a minimal footprint. Their power lies in the shared assumptions between the name’s author and its reader. Whether distinguishing hosts (host1, host2), signaling protocol versions (http2), or denoting matrix dimensions (m33), digits rely on prior knowledge to be effective. This makes them:
  • Easy to understand when used in well-known conventions (e.g., 3D, utf8)
  • Hard to understand when overused without documentation or when the reader lacks background information/experience
Digits are structural shortcuts in the mental models of developers; a quick way to convey a lot of information in a small number of characters.
5.3 Prepositions in Identifiers
Overview.
Prepositions in identifiers express spatial, temporal, or logical relationships. They are the most versatile (i.e., most axial codes) and frequently used closed-class grammatical structure in our dataset. Prepositions typically convey transformation, control conditions, event triggers, source origin, or context membership. Because only a subset of these are dual axis (Boolean Flow), we inline the definitions with our examples, unlike with Digits, where we separated them.
Axial Codes.
Through axial coding, we identified several recurring behavioral roles that prepositions play in identifier names. These axial codes describe the functional semantics conveyed by the preposition within the naming context:
  1. 1.
    Type Casting / Interpretation (38 items)
    Definition: This group captures identifiers that signify transformation from one type, format, or abstraction to another.
    Examples: str_2_int, as_field
    Grammar patterns:
    • P N (18)
    • P NM N (9)
    • N P N (4)
    • V P N (2)
    • P NM NM N (2)
    • P V (1)
    • NM N P N (1)
    • V P (1)
  2. 2.
    Position / Ordering in Time or Space (28 items)
    Definition: This group captures identifiers that indicate relative position or sequencing within a spatial, temporal, or execution context.
    Examples: before_major, after_first_batch
    Grammar patterns:
    • P N (8)
    • P (4)
    • N P N (3)
    • P NM N (3)
    • V P N (3)
    • P V (2)
    • V P (2)
    • NM N P N (2)
    • N P (1)
  3. 3.
    Boolean Flow / Control Flag (26 items)
    Definition: This group captures identifiers that encode boolean flags which both guard execution and describe the behavior they enable. This group is somewhat special, as their name implies other axial codes, but they are boolean variables. Thus, many of the identifiers in this group are dual-axis, where the 1st axis is boolean, and the 2nd is one of the other preposition axes. These variables are typically guards, used in branching logic that:
    • Activate based on position or sequencing (e.g., after_equals)
    • Govern strategy or type casting/interpretation behavior (e.g., for_backprop, as_array)
    • Reflect data provenance or deferred logic (e.g., from_docker_config, wait_for_reload)
    Examples: obsess_over_host, for_backprop
    Grammar patterns:
    • P N (12)
    • P NM N (6)
    • V P N (2)
    • V P (2)
    • N P N (2)
    • N P (1)
    • P (1)
  4. 4.
    Data Source / Origin (20 items)
    Definition: This group captures identifiers that refer to the source from which data or configuration is retrieved.
    Examples: from_context, from_id
    Grammar patterns:
    • P N (10)
    • P NM N (1)
    • N P N (2)
    • P (3)
    • NM N P N (1)
    • V P (1)
    • N P (2)
  5. 5.
    Event Callback / Trigger (17 items)
    Definition: This group captures identifiers that define behavior executed in response to user or system events.
    Examples: on_reason, on_start
    Grammar patterns:
    • P N (6)
    • P NM N (5)
    • P NM NM N (4)
    • V P N (1)
    • NM N P N (1)
  6. 6.
    Deferred Processing / Pending Action (13 items)
    Definition: This group captures identifiers that signal actions or data awaiting future handling.
    Examples: to_ack, to_count
    Grammar patterns:
    • P V (10)
    • P N (2)
    • P NM N (1)
  7. 7.
    Unit-Based Decomposition / Measurement (11 items)
    Definition: This group captures identifiers that describe per-unit measurement, processing, or aggregation.
    Examples: down_time, size_in_datum
    Grammar patterns:
    • NPL P N (8)
    • P N (1)
    • N P N (1)
    • NM N P N (1)
  8. 8.
    Purpose / Role Annotation (10 items)
    Definition: This group captures identifiers that clarify the functional role or use-case of a value.
    Examples: for_avg, for_class
    Grammar patterns:
    • P N (6)
    • NM N P N (2)
    • P NM N (1)
    • V P (1)
  9. 9.
    Data Movement / Transfer (9 items)
    Definition: This group captures identifiers that represent movement of data or control between locations, buffers, or components.
    Examples: to_repo, to_header
    Grammar patterns:
    • P N (3)
    • N P (3)
    • P NM N (1)
    • NM N P N (1)
    • P NM NM N (1)
  10. 10.
    Operation Basis / Strategy (8 items)
    Definition: This group captures identifiers that describe the method, or trait that determines how operations may/should be carried out.
    Examples: extend_by_hexahedron, with_unary_operator
    Grammar patterns:
    • P N (2)
    • P NM N (2)
    • V P N (2)
    • P (1)
    • V P (1)
  11. 11.
    Membership / Peer Grouping (7 items)
    Definition: This group captures identifiers that signal inclusion in a group, scope, or set of peer entities.
    Examples: in_neighbour_heap, in_for
    Grammar patterns:
    • P (2)
    • P N (1)
    • P NM N (1)
    • V P N (1)
    • V P (1)
    • N P (1)
  12. 12.
    Mathematical / Constraint Context (2 items)
    Definition: This group captures identifiers that encode numerical limits, bounds, or ratios that constrain behavior.
    Examples: over_size, vmax_over_base
    Grammar patterns:
    • P N (1)
    • N P N (1)
Selective Coding Insight.
Prepositions in identifier names serve as compact, highly expressive relational markers. Across the dataset, prepositions consistently support four core semantic roles:
  • Transformation and Directionality: Prepositions like to, from, and as signal type casting, movement, or format conversion.
  • Execution and Conditional Control: Prepositions such as after, on, and for often signal when or whether an action should occur, especially within event-driven operations and boolean flags that gate execution.
  • Role and Configuration Semantics: Prepositions like with, by, and in clarify how values contribute to a process or how behavior is scoped or grouped.
  • Quantification and Unit-Based Aggregation: Prepositions such as per and in describe how quantities are measured, normalized, or decomposed across units (e.g., iterations_per_sample, size_in_datum).
  • Future-Intent or Deferred Action: Especially with to, some identifiers encode pending or scheduled behavior (e.g., to_merge, wait_for_reload).
Importantly, boolean flags that include prepositions do not form a distinct behavioral class, but instead overlay these four functions; gating type conversions, controlling source-based logic, or scoping strategies. These flags act as behavioral summaries, where the identifier directly reflects the guarded behavior (e.g., send_to_buffer reflects that the guarded code sends data to a buffer).
In short, prepositions make invisible system relationships visible. They map the logic of control, transformation, and association directly into identifier structure, enabling expressive, intention-revealing naming in complex systems.
5.4 Determiners in Identifiers
Overview.
Determiners in identifiers help interpret values in relation to a set. They often signal positional reasoning, filtering criteria, relative thresholds, control flow, or scoping rules. In our analysis, we treat terms like next and last as determiners, even though they are typically categorized as adjectives in general English. In source code, however, these terms function more like determiners because they specify a particular entity within a sequence or collection rather than merely describing its properties. For example, the next pointer in a linked list does not describe a type of pointer, but rather identifies the specific node that follows in the structure. In this way, such terms serve a determinative function.
Axial Codes.
We identified the following eight categories of determiner-based behavior:
  1. 1.
    Temporal / Most Recent Element (60 items)
    Definition: This group captures identifiers that refer to the most recently computed, stored, or observed value, often used for computing prior state, and in sequence-based data structures.
    Examples: last_bucket, last_builder
    Grammar patterns:
    • DT N (32)
    • DT NM N (19)
    • DT NM NM N (4)
    • DT V (2)
    • V DT N (2)
    • DT NPL (1)
  2. 2.
    Temporal / Upcoming Element (54 items)
    Definition: This group captures identifiers that denote the next item in a sequence or timeline, often used in look-ahead and sequence-based data structures.
    Examples: next_tex, next_bar
    Grammar patterns:
    • DT N (35)
    • DT NM N (9)
    • DT V (3)
    • N DT (3)
    • DT NPL (2)
    • DT NM NM N (1)
    • V DT N (1)
  3. 3.
    Population / Subpopulation Reference (42 items)
    Definition: This group captures identifiers that reference a population or subset, typically using quantifiers like all, any, or some to guide iteration, filtering, or policy logic.
    Examples: any_diffuse, all_set
    Grammar patterns:
    • DT NPL (13)
    • DT N (9)
    • V DT (6)
    • DT NM NPL (4)
    • V DT NPL (4)
    • DT NM N (2)
    • N DT (2)
    • DT V (1)
    • V DT N (1)
  4. 4.
    Immediate Context Reference (26 items)
    Definition: This group captures identifiers that refer to the current instance, scope, or runtime context—emphasizing locality, such as this, another, or a.
    Examples: this_node, another_id
    Grammar patterns:
    • DT N (17)
    • DT NM N (6)
    • DT NM NM N (1)
    • N DT (1)
    • V DT N (1)
  5. 5.
    Negation / Exclusion Flag (18 items)
    Definition: This group captures identifiers that indicate something is explicitly disabled, excluded, or absent; commonly using no to toggle features or signal null conditions.
    Examples: no_callback, no_log
    Grammar patterns:
    • DT N (12)
    • DT NM N (2)
    • DT NPL (2)
    • DT NM NPL (1)
    • DT V (1)
  6. 6.
    Quantity Threshold / Optional Extensibility (4 items)
    Definition: This group captures identifiers that express minimum thresholds, or the possibility of extending beyond a baseline.
    Examples: enough_memory, more_data
    Grammar patterns:
    • DT N (2)
    • DT NPL (2)
  7. 7.
    Default / Fallback Value Representation (2 items)
    Definition: This group captures identifiers that represent placeholder or fallback values, used when a field must be filled or a default condition must be satisfied.
    Examples: a_void, no_val
    Grammar patterns:
    • DT N (2)
  8. 8.
    Boolean Multi-Condition Test (2 items)
    Definition: This group captures boolean identifiers representing conjunctions of multiple conditions, usually requiring all to be satisfied (e.g., both X and Y must be true).
    Examples: both_empty_selection, both_NonEmpty_Selection
    Grammar patterns:
    • DT NM N (2)
Selective Coding Insight.
Determiner-based identifiers help interpret values in relation to a set—by signaling position, filtering criteria, thresholds, or scoping rules. These are closed-category terms that enable programmers to express set logic, entity selection, and relative capacity or validity. They typically support:
  • Positional reasoning (next, last, this): Indicates where a value occurs in a temporal or structural sequence, helping to track state progression, history, or future execution.
  • Population membership and filtering (some, any, each, least, which): Refers to selecting or referencing members within a larger set, expressing scope, quantification, or comparison.
  • Thresholding and extensibility (enough, more, additional): Indicates whether a minimum condition is met or whether more values can be included beyond a base requirement.
  • Identity negation or fallback (no, none, a, without): Flags exclusion, absence, or placeholder values—often tied to feature toggles or default logic.
5.5 Conjunctions in Identifiers
Overview.
Conjunction-based identifiers are rare but expressive. They signal compound behavior, dual-mode interfaces, or gated logic—often making hidden control flow or semantic relationships visible. Their rarity likely stems from the fact that developers often express conjunctions in logic rather than names. But when used, they highlight either an intent to foreground behavior or to capture structural duality within a single name.
Axial Codes.
We identified seven categories of conjunctional behavior, each reflecting a different type of pairing, conditionality, or combination:
  1. 1.
    Data Pair / Composite Value (7 items)
    Definition: This group captures identifiers that hold or refer to two values used together or in alternation, typically for a shared behavioral role or composite purpose.
    Examples: data_or_diff, function_and_data
    Grammar patterns:
    • N CJ N (6)
    • V CJ N (1)
  2. 2.
    Guarded Action / Conditional Enablement (6 items)
    Definition: This group captures identifiers that encode actions gated by internal logic; executing only if a condition is satisfied. The conjunction expresses conditional enablement or guarded behavior.
    Examples: if_present, if_unique
    Grammar patterns:
    • CJ NM (2)
    • V CJ N (2)
    • V CJ V (1)
    • V CJ VM P (1)
  3. 3.
    Combined Action / Sequential Behavior (3 items)
    Definition: This group captures identifiers that describe a sequence of operations performed together, often representing merged behaviors.
    Examples: hash_and_save, print_and_free_json
    Grammar patterns:
    • V CJ V (1)
    • V CJ V N (1)
    • V N CJ N (1)
  4. 4.
    Shared Interface for Alternatives (1 item)
    Definition: This group captures identifiers that define a shared interface or behavior over mutually exclusive alternatives, with the conjunction indicating a choice, not a combination.
    Example: generate_key_or_iv
    Grammar pattern:
    • V N CJ N (1)
  5. 5.
    Combined Configuration / UI Concept (1 item)
    Definition: This group captures identifiers that refer to compound interface or configuration concepts, often blending multiple traits into a unified design or behavioral setting.
    Example: look_and_feel
    Grammar pattern:
    • NM CJ NM (1)
  6. 6.
    Boolean Concept Name (1 item)
    Definition: This group captures identifiers that encode a named logical or boolean relationship, usually by treating the conjunction itself as a symbolic concept.
    Example: and
    Grammar pattern:
    • CJ (1)
  7. 7.
    Boolean Multi-Condition Test (1 item)
    Definition: This group captures identifiers that evaluate multiple conditions simultaneously; typically for readiness or validation checks, returning true only if all constraints are met.
    Example: null_or_empty
    Grammar pattern:
    • NM CJ NM (1)
Selective Coding Insight.
Conjunction-based identifiers are especially useful when modeling:
  • Duality: Representing more than one entity or mode simultaneously (e.g., input_and_output, key_or_iv)
  • Mutual Exclusion: Encoding choices between alternatives—only one active at a time (e.g., stream_or_cache)
  • Preconditions: Embedding logic into the name that would otherwise be hidden in branching statements (e.g., load_if_needed, trigger_if_active)
Conjunctions are the rarest category in our data, and while it is difficult to draw firm conclusions about them, it is clear that ‘and’, ‘or’, and ‘if’ are go-to conjunctions, particularly for Data Pairs and Guarded actions.
5.6 Cross-Category Synthesis
Across digits, determiners, prepositions, and conjunctions, developers use closed-class grammatical structures to encode compact, behavior-rich semantics in identifiers. While each part-of-speech (POS) category exhibits distinct tendencies, analysis of grammar patterns reveals broader functional themes and stylistic consistencies across categories.
Boolean Semantics and Execution Control.
Our first cross-category behavior is the use of closed-class elements to encode boolean conditions, execution control, or logical gating:
  1. 1.
    Determiners such as no, some, this, and both signal presence, exclusion, or multi-condition boolean evaluation.
  2. 2.
    When used as booleans, Prepositions like as, and with tend to guard sections of code that implement the behavior described in the identifier name.
  3. 3.
    Conjunctions surface explicitly in guarded or compound logic names (e.g., load_if_enabled, both_ready) using patterns like V CJ N, NM CJ NM.
It is interesting that booleans appear in all three of these contexts, but each is a different flavor; a way of expressing behavior that is unique to the closed-category terms used in the boolean identifier.
Control Flow and Event Signaling Across Categories.
Closed-category terms across all four categories reflect a tendency to encode temporal, reactive, or preconditioned behavior:
  1. 1.
    Prepositions like on, before, after, and by appear in structures such as P N and V P N, signaling timing, triggers, or basis for operation.
  2. 2.
    Conjunctions explicitly model control conditions (if, and) or mutual exclusivity (or), often appearing in V CJ N or N CJ N structures.
  3. 3.
    Determiners frequently encode sequence through next and last, realized in DT N and DT NM N patterns.
  4. 4.
    Digits imply procedural differentiation (method1, step2) or timeline indexing when appearing in coordinated identifiers (m31, m32).
These names act as micro-control structures, embedding state transitions and flow logic directly into identifier names to help the reader understand when or how an identifier will/should be used.
Multi-Dimensional Semantic Layering.
Grammar pattern analysis highlights how identifiers stack multiple behavioral dimensions:
  1. 1.
    Prepositions convey direction, transformation, measurement, and order
  2. 2.
    Determiners convey selection, quantity, and scope
  3. 3.
    Digits embed indexing, uniqueness, and domain roles
  4. 4.
    Conjunctions encode logic composition and structural alternatives.
These layered forms operate as semantic shortcuts to express complex behavior with very few words. They compress conditions, transformations, order, and relationships into concise forms.
Finally, the grammar patterns observed across our axial codings provide structural insight into how behavioral semantics are composed. When the closed-category term appears as the first token in a grammar pattern, such as in DT NM N or P NM N, it typically modifies or qualifies a single operand, forming a unary relation (e.g., temporal status or transformation of a noun phrase). In contrast, when the closed term is flanked by open-class terms, such as in N P N or N CJ N, the structure reflects a binary relation: two operands connected through a behavioral or logical relationship (e.g., data flow or choice).
By combining our axial and selective codes with these syntactic patterns, we gain a fuller picture of identifier meaning: the open-category terms indicate *which* entities are involved, while the closed-category term signals *how* they are related or behave with respect to one another.
5.7 Summary of RQ1
Through grounded theory analysis of closed-category terms in identifiers, we have uncovered and explored the ways in which these compact grammatical forms play a central role in expressing program behavior. Each part-of-speech category contributes distinct semantic functions, ranging from transformation and scoping to control flow and logical composition.
Together, they reveal how developers construct concise, behavior-rich identifiers that encode structure, timing, intent, and logic. Whether signaling preconditions (load_if_enabled), alternatives (data_or_diff), state (last_bucket), or structural roles (col1), these terms form a functional lexicon that bridges source code, cognition, and context.
Table 8: Top 10 terms per closed-category part-of-speech tag
Determiner Conjunction Digit Preposition
last (79, 25.65%) and (18, 36.00%) 1 (77, 27.21%) to (76, 19.10%)
next (69, 22.40%) if (16, 32.00%) 2 (71, 25.09%) for (37, 9.30%)
all (52, 16.88%) or (13, 26.00%) 0 (20, 7.07%) on (36, 9.05%)
no (31, 10.06%) than (1, 2.00%) 3 (17, 6.01%) as (35, 8.79%)
this (21, 6.82%) since (1, 2.00%) 4 (15, 5.30%) from (33, 8.29%)
each (7, 2.27%) when (1, 2.00%) 16 (13, 4.59%) in (25, 6.28%)
the (6, 1.95%) - 8 (8, 2.83%) after (21, 5.28%)
more (5, 1.62%) - 6 (6, 2.12%) 2 (to) (17, 4.27%)
a (4, 1.30%) - 64 (4, 1.41%) of (15, 3.77%)
some (4, 1.30%) - 32 (3, 1.06%) with (14, 3.52%)
6 Evaluation of RQ2: How do closed-category terms correlate with structural, programming language, and domain-specific contexts in software?
One interesting aspect of closed-category terms is that they appear in different contexts within source code with varying frequency. This variation provides insight into how developers use these terms to express different types of meaning. For RQ2, we investigate how closed-category terms correlate with three types of context: (1) the local programming context in which a variable is declared (e.g., Function, Attribute), (2) the programming language of the source code in which the identifier was found, and (3) the broader system-level domain of the software in which it appears (e.g., domain-specific vs general-purpose projects). This 3-way perspective allows us to examine both how these terms are used within individual source code structures, between programming languages, and how they reflect distinctions across different kinds of systems.
We begin by analyzing the distribution of four closed categories: prepositions, determiners, conjunctions, and digits, across five programming contexts and three programming languages. We discuss which categories are most frequent in which contexts/languages and consider how those patterns may reflect the communicative goals of the developer. We then extend this analysis to system-level domain context, comparing the normalized frequency of closed-category term usage between domain-specific and general-purpose systems.
6.1 Closed-Category Term Usage Across Programming Contexts, Programming Languages, and System Domains
We now examine how differing contexts and closed-category grammar patterns relate to one another, and whether programming language further conditions their usage. We begin by analyzing cross-language correlations in closed-category term usage, followed by an exploration of correlations in how closed-category terms are used in different program contexts. We provide Table 8, which shows frequencies and percentages for PoS and terms, to help the reader understand what types of terms are most prevalent. However, for this RQ we rely primarily on Tables 9, 10, 11, and 12, which show the results of our Pearson Chi-squared tests and Standardized Pearson residuals. Using these, we highlight common patterns, terms, and the contexts or languages to which these patterns are correlated.
6.1.1 Language-Specific Differences in Closed-Category Term Usage
Starting with an analysis of closed category terms and programming language, our null hypothesis is that there is no relationship between identifiers containing closed category terms and the programming language in which they appear. Our alternative hypothesis is that there is a relationship between identifiers that contain closed category terms and the programming language.
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We count how many times each closed-category PoS appears in C++, Java, or C code by analyzing all 1,001 identifiers that contain closed-category terms. For example, we might find that there were 20 Digits in our data set found in C++ code, and 5 Digits in Java code. Once we have these frequencies, we apply the Chi-Square test, and Standardized Pearson residuals w/Bonferonni correction to determine overall significance, and per-part-of-speech significance, respectively.
Table 9: Results of Pearson’s Chi-Squared Test. df = 6, $\alpha$ = 0.05, critical value = 12.592, test statistic = 4.291
C C++ Java Chi-square per row
D 0.451887 0.275300 1.282415 2.009603
DT 0.457607 0.036988 0.678404 1.172999
P 0.056554 0.015121 0.116729 0.188405
CJ 0.621604 0.157629 0.140876 0.920108
Chi-square per column 1.587652 0.485038 2.218424 4.291115
Table 10: Standardized Pearson Residuals Results. With Bonferroni Correction, a significant result is $\alpha$ = 0.05/12 = 0.0042, which translates to a ±2.87 critical value
C C++ Java
D 0.953270 0.740778 -1.649081
DT -0.986465 -0.279221 1.233406
P -0.367728 -0.189311 0.542514
CJ 0.983056 -0.492859 -0.480581
Results. The Chi-square test for programming language (Table 9) did not produce a statistically significant result. Thus, we do not reject the null hypothesis: there is no strong evidence that closed-category tag usage differs significantly by programming language. However, exploratory analysis of the Standardized Pearson residuals in Table 10 offers insight into modest trends worth noting:
  • Digits (D) are modestly underrepresented in Java (residual = –1.65), suggesting a mild tendency to avoid numeric suffixes in Java naming.
  • Determiners (DT) are slightly overrepresented in Java (residual = 1.23), potentially reflecting more frequent use of quantifying or contextual modifiers.
Summary.
While we did not find significant statistical evidence linking closed-category tag usage to programming language, the residual analysis and qualitative trends suggest mild idiomatic differences, particularly around digit usage and determiner phrasing. For example, in our Axial Code data from RQ1, Population/Subpopulation Reference identifiers were found in Java (21, 50%) and C++ (18, 42%) more than in C(3, 7%). These patterns may reflect broader stylistic conventions or design idioms of each language, but should be interpreted cautiously given the statistical outcome.
Table 11: Results of Pearson’s Chi-Squared Test. df = 12, $\alpha=0.05$, critical value = 21.026, test statistic = 88.893567.
ATTRIBUTE CLASS DECLARATION FUNCTION PARAMETER Chi-square per row
D 0.44932 16.031139 0.776138 15.916173 10.634467 43.807237
DT 0.511266 6.398601 2.14863 0.959251 0.171805 10.189553
P 0.332388 0.506133 3.498335 8.724009 3.63817 16.699033
CJ 3.366551 1.027972 0.233783 12.071429 1.498009 18.197744
Chi-square per column 4.659525 23.963845 6.656886 37.670861 15.942451 88.893567
Table 12: Standardized Pearson Residuals. With Bonferroni correction, a significant result is $\alpha=0.05/20=0.0025$, which translates to a ±$\pm 3.02$ critical value.
ATTRIBUTE CLASS DECLARATION FUNCTION PARAMETER
D -0.905532 4.719156 1.1768 -5.505053 4.254103
DT 0.993307 -3.065907 2.013483 -1.389769 -0.556035
P 0.849262 -0.91434 -2.724314 4.444202 -2.713221
CJ -2.179409 -1.050735 -0.567884 4.21543 -1.403873
6.1.2 Context-Specific Differences in Closed-Category Term Usage
Next, we analyze the correlation between closed-category terms and program contexts such as Function names, Attributes, and Class names. Our null hypothesis is that there is no relationship between identifiers containing closed-category terms and the context in which they appear. Our alternative hypothesis is that there is a relationship between identifiers containing closed-category terms and the context in which they appear.
Methodology. To perform our Chi-Square test, we use the CCID described in Section 4.1. We analyzed all 1,001 identifiers that contained closed-category terms, and count how many times a closed-category term appears in one of our five code contexts: Attribute, Function, Class, Declaration, or Parameter. Once we have these frequencies, we apply the Chi-Square test, and Standardized Pearson residuals w/Bonferonni correction to determine overall significance, and per-part-of-speech significance, respectively.
Results. The Chi-squared test for context (Table 11) shows a significant result (88.89 $>$ 21.026), allowing us to reject the null hypothesis. As before, we analyze the Standardized Pearson Residuals (Table 12) to understand where the largest deviations appeared.
Conjunctions (CJ). Closed-category grammar patterns that include conjunctions typically feature the terms ‘and’, ‘or’, or ‘if’, as reflected in Table 8. Although rare overall, these patterns are significantly positively correlated with function names (Standardized Pearson residual = 4.22, Table 12). This indicates that when conjunctions do appear, they are far more likely to occur in function names than in other contexts.
The selective coding data from RQ1 offers an explanation for this pattern. Conjunction-based grammar patterns tend to express compound logic, dual-purpose behavior, or guarded activation, which are most relevant when naming behaviors or actions rather than static values. For example, Guarded Action / Conditional Enablement patterns such as load_if_needed or activate_if_enabled appear in function names to encode preconditions or gating logic directly into the identifier.
Conjunction-based names are largely absent from declarations and classes, likely because those contexts do not typically represent conditional or compound operations. The data supports the interpretation that developers strategically use conjunctions in function names to foreground complex control logic or behavioral nuance at the point of execution.
While we did find a significant correlation with Conjunctions and functions, it is important to recall that we have a limited number of Conjunctions in our dataset, at 16, meaning that while we have found potential trends, further research is required.
Determiners (DT). Closed-category grammar patterns that include determiners typically feature the terms last, next, all, no, or this, as shown in Table 8. While determiners are not significantly positively correlated with any specific context, they are modestly negatively correlated with class names (Standardized Pearson residual = –3.07, Table 12), suggesting that developers tend to avoid determiner-based grammar patterns in class names.
The selective coding analysis offers a plausible explanation: the most common roles for determiners involve expressing temporal or positional relationships—such as Temporal / Most Recent Element and Temporal / Upcoming Element (over 100 instances)—as well as set-based semantics, such as Population / Subpopulation Reference (38 instances). These patterns commonly use terms like next, last, prev, all, and any to indicate an element’s position in a sequence or its membership in a filtered subset.
These naming strategies are well-suited to attributes, parameters, and declarations, where variables often represent dynamic state or bounded subsets. In contrast, class names are generally used to describe abstract data types or roles, where positional or filtering semantics are less relevant. The relative absence of determiners in class contexts thus reflects their semantic focus: determiners foreground state, scope, or specificity, whereas class names typically signal structural purpose or generalization.
Digits (D). Closed-category grammar patterns that include digits often feature numerals such as 1, 2, 0, 3, and 4, as shown in Table 8. Digits are significantly positively correlated with parameter names and class names, and significantly negatively correlated with function names (Standardized Pearson residual = 4.72 for class names, 4.25 for parameters, and –5.51 for functions; Table 12). Notably, digits are the only closed-category category to exhibit a positive correlation with class names.
The selective coding data sheds light on this trend. The most frequent digit-related patterns in our dataset fall under Distinguisher × Human-Named Convention and Distinguisher × Locally Specific Concept. These naming strategies are used to distinguish among similar entities (e.g., arg1, arg2, tile3) or to embed system-specific references (e.g., m34, cp437) into variable or type names. Such distinctions are especially useful in parameters and declarations, where there is no syntactic support for disambiguation outside of naming.
In contrast, function-level disambiguation is often handled by the language itself, through overloading, polymorphism, or naming conventions focused on behavior; making digits largely unnecessary or even undesirable in that context. Their absence from function names reflects this shift: digits encode identity, that is, they act as traceability to specific domain concepts, and distinguishing entities with similar names, rather than encoding purpose or behavior.
Taken together, these findings suggest that digits serve primarily as disambiguators or protocol markers rather than communicative devices for expressing behavior. Their presence in class and parameter names signals static or structural variation, while their avoidance in function names underscores a developer preference for meaningful, descriptive action labels over numerical markers.
Prepositions (P). Closed-category grammar patterns that include prepositions frequently feature terms such as to, for, as, on, or from, as shown in Table 8. These patterns are significantly positively correlated with function names (Standardized Pearson residual = 4.44, Table 12), suggesting that developers are particularly likely to use prepositional grammar when naming behaviors or operations.
This strong correlation reflects the behavioral semantics that prepositions convey in identifier names. As detailed in our selective coding, prepositions frequently express directionality, transformation, conditional activation, or event-driven execution; all of which are very function-oriented behaviors, requiring action to be taken. Overall, prepositions help scope, qualify, and clarify a function’s behavior. Their strong correlation with function names supports the interpretation that developers use them deliberately to encode operational semantics directly into the name, especially in contexts involving transformation, control flow, or event handling.
Summary.
The results of our analysis support the alternative hypothesis: closed-category parts of speech are meaningfully correlated with specific roles and contexts in source code. Prepositions and conjunctions appear more frequently in function names, where they help express behavioral nuances such as guarded actions, type casting, or alternative execution paths. Digits, by contrast, are most commonly found in class names and parameter declarations, where they signal disambiguation, indexing, or versioning; identifiers rooted in identity rather than behavior or purpose.
Figure 2: Global Mann-Whitney U test significance across thresholds, showing divergence between domain-specific and general systems. Peaks at 0.6 and 0.8 suggest the importance of both ubiquitous and moderately specific closed-category terms.
Figure 3: Per-category Mann-Whitney U test significance across thresholds. Prepositions dominate across thresholds, while conjunctions and digits contribute more variably.
Figure 4: Cliff’s Delta for closed-category terms across system support thresholds.
Table 13: Systems and System Domains Selected Based on Axial Codes for each Closed-Category Type
Closed-Category Type Axial Code Domain GitHub URL Language Commit Hash
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/msgpack/msgpack-c/ C 306d59d52f752388f276aa5244cdd56be3ad75e0
Prepositions Type Casting / Interpretation Serialization/Deserialization Libraries https://github.com/open-source-parsers/jsoncpp.git C++ ca98c98457b1163cca1f7d8db62827c115fec6d1
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/swig/swig C++ 4167b0275adc574f86914e42fb5ef9639b2b25a5
Prepositions Type Casting / Interpretation Polyglot Interop Tools / Type Bridge Layers https://github.com/pybind/pybind11 C++ bc4a66dff0464d0c87291b00a3688999fea5bedc
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/boostorg/container C++ b13c1fe8a4921328c9ecdbdfc90431d73b2060ac
Prepositions Position / Ordering in Time or Space Data Structure / Algorithm Libraries https://github.com/apache/commons-collections Java a3dca96b37f728aa191ca4b077b6604f2e3d2daa
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/TinyCC/tinycc C 36ff4f52b518e1043b3aa2f0a466392c08b33978
Prepositions Position / Ordering in Time or Space Compiler / Intermediate Representation Tools https://github.com/rose-compiler/rose C 5c9d8859d38a320d577d7f592a98f4fb762051fc
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/quartz-scheduler/quartz Java df059e2a6c22ad1466476ae246025f31caffa6e5
Determiners Position in Sequence (Temporal / Recent & Upcoming) Job Queues / Schedulers https://github.com/PerMalmberg/libcron C++ aa3d4368d59abb9d51199755df7419194bc3e42e
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/antlr/antlr4 Java cb850120378d000e97da38afb852aa3c554b81f5
Determiners Position in Sequence (Temporal / Recent & Upcoming) Parser Generators / Token Stream Libraries https://github.com/akimd/bison C 25b3d0e1a3f97a33615099e4b211f3953990c203
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/apache/arrow C++ 4d566e60c5a12f047e1c0f36b9733a56513a7d43
Determiners Population / Subpopulation Reference Dataframe / Matrix Libraries https://github.com/jtablesaw/tablesaw Java 5b27ac45e00396fd6dc30572da48ba5181d81ee2
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/pjreddie/darknet C f6afaabcdf85f77e7aff2ec55c020c0e297c77f9
Determiners Population / Subpopulation Reference ML Preprocessing / Feature Selection https://github.com/haifengl/smile Java d18696f3f3bb198a8ee4a3c7485942508e59080a
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/spring-projects/spring-boot Java c6045c3111c43bd7b0f99e6c2858bfb2999e358f
Conjunctions Guarded Action / Conditional Enablement Feature Flag Systems / Config-Driven Execution https://github.com/lightbend/config Java d279841c425e70645feacbdc03e41b3b022ad1f9
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/GNOME/gtk C 17bbeb4bc5556a897198613f872543a9a1b4a240
Conjunctions Guarded Action / Conditional Enablement UI Libraries / Event Dispatch Systems https://github.com/ocornut/imgui C++ 75964a9860d497ce3a4b0c482a093e664971466c
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/openssl/openssl C 29864f2b0f1046177e8048a5b17440893d3f9425
Conjunctions Data Pair / Composite Value Cryptographic Libraries https://github.com/jedisct1/libsodium C 72fdf1b86c476e81a3352292962423e76cb141d8
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/apache/parquet-java Java c780ffa09be559ca55feb1c0ba996554dd416dea
Conjunctions Data Pair / Composite Value Multi-format I/O Libraries https://github.com/libjpeg-turbo/libjpeg-turbo C 2a0c86278249e7a3c3429caff24c06a50048d772
Digits Distinguisher × Human-Named Convention GUI Builders / Form Designers https://github.com/qt/qtbase C++ 9a06c8ae8c0c13d9a84ee18768e1c68169c31043
Digits Distinguisher × Human-Named Convention Code Generators / Macro Frameworks https://github.com/jhipster/jhipster-bom Java 82e1ab06812348569b80711fc80568575e533c8e
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/godotengine/godot C++ d00f25ecd92f100aedbab4ef4622b95900e9cf01
Digits Distinguisher × Locally Specific Concept Game Engines / Grid-based Games https://github.com/jMonkeyEngine/jmonkeyengine Java dcb22624b426c46805c1468e90adcd0365cd3083
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/PX4/eigen C++ 7cf1c0179eb0f5499dfc1bffbd229783a7865fe1
Digits Distinguisher × Locally Specific Concept Scientific Computing / Matrix Libraries https://github.com/OpenMathLib/OpenBLAS C 70dff3b84fff151f3e58bf54ce54a5e763f29cac
6.1.3 Closed-Category Term Usage Across System Domains
Having established correlations between closed-category terms, source code context, and programming language, we now turn to a broader question: do these terms also vary with the domain of the software system itself? This sub-question allows us to further test our central hypothesis, that closed-category terms are not used arbitrarily, but instead reflect domain-relevant distinctions in how behavior and structure are communicated. If certain domains make more frequent or specialized use of closed-category terms, this suggests that such terms play a role in expressing concepts tightly coupled to those domains. Understanding and appropriately using these terms may therefore be critical for accurate communication of behavior in domain-specific software.
Methodology. In RQ1, we developed a set of Axial Codes to describe the behavioral roles of closed-category terms in identifiers. To explore their importance at the level of system domain, we selected the two most common Axial Codes from each closed-category group (e.g., Prepositions, Determiners). For each code, we identified two software domains that we hypothesized would frequently use identifiers expressing that behavior. For example, in the Preposition group, the top two Axial Codes were:
  • Type Casting / Interpretation
  • Position / Ordering in Time / Space / Execution Context
Based on these, we selected four relevant software domains:
  • For Type Casting / Interpretation:
    • Serialization/deserialization libraries
    • Polyglot interop tools or type bridge layers
  • For Position / Ordering in Time / Space / Execution Context:
    • Data structure and algorithm libraries
    • Compiler or intermediate representation (IR) tooling
Table 13 lists all selected systems, the domain they represent, and the Axial Code that motivated their inclusion. To fit the table, we omitted a few details like system size; this can be found in our open data set (Section 11). We analyzed identifiers drawn from five programming contexts—attributes, parameters, functions, declarations, and class names—across two groups of systems: one curated for domain-specific relevance and one composed of general-purpose projects, which were used to construct the CCID (Table 3). The general-purpose group serves as a baseline, as these systems were not selected based on any particular domain hypothesis. Our underlying assumption is that, if closed-category terms are meaningfully correlated with domain-specific concerns, we will observe statistically significant differences in their usage between these two groups.
For each system, we extracted all identifiers and segmented them using Spiral HuckaSpiral . We then filtered out all terms that are neither digits nor included in our predefined lists of closed-category terms (as defined in Section4). After filtering, we compute the normalized frequency of closed-category term usage by dividing the count of qualifying terms by the system’s total lines of code. To assess whether differences in usage were statistically meaningful, we apply a Mann-Whitney U test to compare the distributions between domain-specific and general-purpose systems.
Results. To mitigate the risk of a small number of systems dominating the term distribution, and to better understand how widely closed-category terms are used, we introduce a support threshold that controls how many systems a term must appear in to be included in the Mann-Whitney U test. Increasing the threshold emphasizes more widely used (ubiquitous) closed-category terms; decreasing it emphasizes more narrowly distributed (specific) closed-category terms that may signal domain-specific behavior. Significance at high thresholds implies that there are terms that are important to all of our domain-specific systems; significance at lower thresholds implies that there are subsets of the domain-systems that make use of terms that are not very universal, but nevertheless set these systems apart from the general set.
To explore how these different usage profiles affect our results, we conducted a threshold sweep. At each level, a term had to appear in at least a given proportion of systems to be retained. This allowed us to systematically vary our emphasis between ubiquity (terms common across many systems) and specificity (terms concentrated in a smaller, domain-aligned subset). The results, shown in Figure 2, reveal the strongest distributional divergence at thresholds around 0.6 and 0.8. These peaks suggest that both common and moderately specific terms help distinguish domain-specific systems. By contrast, thresholds between 0.1 and 0.3 yielded little significance, likely reflecting linguistic noise from terms with low usage or ambiguous semantic function.
We repeat the analysis at the level of individual closed-category types (prepositions, determiners, conjunctions, digits) to identify which groups drive the observed differences. As shown in Figure 3, prepositions exhibit consistently strong significance across thresholds, particularly between 0.6 and 0.8. Digits and conjunctions show more variable but still notable divergence, while determiners contribute the weakest and least consistent signal. These trends suggest that domain-specific systems rely more heavily on certain linguistic forms, especially prepositions, to express structural or behavioral distinctions central to their design.
To complement the significance testing, we examine Cliff’s Delta as a non-parametric effect size estimate, plotted in Figure 4. This allows us to assess not only whether closed-category usage differs between system types, but also how strongly. The results show that prepositions and digits increasingly favor domain-specific systems at higher thresholds, reaching small to medium effect sizes. Determiners, by contrast, exhibit weak or even negative effect sizes, suggesting a more general-purpose usage profile. Conjunctions remain close to the negligible–small range, with mild domain skew. These patterns reinforce the idea that domain-specific systems do not just differ in which closed-category terms they use, but in how salient those terms are among their most widely reused identifiers.
Summary.
Our findings suggest that domain-specific systems tend to use closed-category terms more frequently than general-purpose baselines, particularly in ways that align with the communicative roles captured by our Selective Codes. While we rely on predefined lists of closed-category terms—without verifying each term’s function in context—our goal in this evaluation was not to establish definitive usage, but to assess whether these terms might play a heightened role in domain-specific software. The statistical results support that possibility. As such, we argue that further research into how closed-category terms contribute to domain-specific expression is both warranted and promising. These findings offer initial evidence that supporting developers in the effective use of such terms could benefit certain styles or domains of software development.
6.2 Summary of RQ2
For RQ2, we examine how closed-category terms correlate with multiple forms of context: (1) source-code-local structure, (2) programming language, and (3) broader system domain. Our findings reveal several consistent trends. First, there is no statistically significant difference in the distribution of closed-category terms across the three programming languages under study, though there are some trends that indicate how they may differ in minor (i.e., non-statistically-significant) ways. Second, source code context plays a significant role: Prepositions and conjunctions are used disproportionately in function names; Digits are significantly positively correlated with parameters and class names while significantly negatively correlated with function names; and Determiners are significantly negatively correlated with class names. These patterns align with the communicative roles uncovered in our Selective Codes, such as the use of prepositions to express behavior or data flow, and digits to distinguish instances or versions.
Finally, we found statistical evidence that domain-specific systems use closed-category terms more frequently than general-purpose ones. This suggests that these terms serve as meaningful signals of domain-relevant behavior. Taken together, our results demonstrate that closed-category terms have specific, purposeful usage in software development.
One of the broader aims of RQ2 is to assess whether closed-category terms are meaningful enough to warrant dedicated study. We argue that their statistically significant correlations with specific code contexts support this aim: such terms appear deliberately and consistently in ways that reflect their natural language functions. While our domain-level comparison relies on predefined lists of closed-category terms, without manual verification of each term’s grammatical role, the results nonetheless suggest that these terms may hold particular communicative importance in domain-specific software. This exploratory finding reinforces the potential value of further study. Supporting the appropriate use of closed-category terms through tools, naming conventions, or educational interventions may ultimately benefit program comprehension and internal quality, particularly in domains where such terms help convey behavioral intent.
7 Related Work
While there are numerous studies on identifier names, this paper represents one of the few to address closed-category terms, and the only paper to do an in-depth study of their usage in open source systems. We discuss relevant related literature below, and how our work can be improved by, or improve upon, their outcomes.
7.1 Part of Speech Taggers
POSSE Gupta:2013 and SWUM HillSWUM:2010 , and SCANL tagger newman_ensemble are part-of-speech taggers created specifically to be run on software identifiers; they are trained to deal with the specialized context in which identifiers appear. Both POSSE and SWUM take advantage of static analysis to provide annotations. For example, they will look at the return type of a function to determine whether the word set is a noun or a verb. Additionally, they are both aware of common naming structures in identifier names. For example, methods are more likely to contain a verb in certain positions within their name (e.g., at the beginning) Gupta:2013 ; HillSWUM:2010 . They leverage this information to help determine what POS to assign different words. Newman et al. Newman_GP compared these taggers on a larger dataset than their original evaluation (1,335 identifiers) using five identifier categories: function, class, attribute, parameter, and declaration statement. They found that SWUM was the most accurate overall, with an average accuracy around 59.4% at the identifier level. Later, Newman et al. created a new tagger that ensembled SWUM, POSSE, and Stanford together, then compared with SWUM, POSSE, and StanfordToutanova:StanfordTagger individually, finding that the ensembled tagger exceeded the others’ performance metrics on identifiers newman_ensemble .
7.2 Human-subjects studies
Several studies use human subjects to understand the influence and importance of different characteristics of identifiers. Our work is largely complementary to these studies, as it can be used in conjunction with data from these studies to create/support naming techniques. Reem et al. reem:2021 did a survey of 1100 professional developers, shedding some light on developer preferences and practices with respect to the content of identifier names, including things such as the use of abbreviations and preferred identifier length. Feitelson et al. feitelson:2022 studied how the information content of identifiers named affected their memorability, and concluded that short names that contain focused information are likely optimal. Felienne et al. Felienne:2024 find, among other things, that while instructors agree on the importance of naming, there is disagreement between their teaching practices. Even internally, teachers are generally inconsistent in how they teach and practice identifier naming in the classroom. The results of their study highlight the importance of increasing our formal understanding of naming, which can help increase and support consistency of teaching materials and practices in the classroom.
7.3 Rename Analysis
Arnoudova et al. Arnaoudova:2014 present an approach to analyze and classify identifier renamings. The authors show the impact of proper naming on minimizing software development effort and find that 68% of developers think recommending identifier names would be useful. They also defined a catalog of linguistic anti-patterns Arnaoudova:2013 . Liu et al.liu2015identifying proposed an approach that recommends a batch of rename operations to code elements closely related to the rename. They also studied the relationship between argument and parameter names to detect naming anomalies and suggest renames liu2016nomen . Peruma et al. Peruma:2018:EIW:3242163.3242169 studied how terms in an identifier change and contextualized these changes by analyzing commit messages using a topic modeler. They later extend this work to include refactorings Perumascam and data type changes PERUMA2020110704 that co-occur with renames. Osumi et al. shinpei:2022 studied terms that were co-renamed with a goal of supporting developers in deciding when identifiers should be renamed together. In particular, they studied how location, data dependencies, type relationships, and inflections affected co-renaming.
These techniques are concerned with examining the structure and semantics of names as they evolve through renames. By contrast, we present the structure and semantics of names as they stand at a single point in the version history of a set of systems. Rename analysis and our work are complementary; our analysis of naming structure can be used to help improve how these techniques analyze changes between two versions of a name by examining changes in their grammar pattern. In particular, since we specifically study closed-category terms, rename analysis can leverage our results to improve its behavior on identifiers that contain these terms. For example, they might use our results to determine when to recommend a closed-category term during a rename operation.
7.4 Identifier Type and Name Generation
There are many recent approaches to appraising identifier names for variables, functions, and classes. Kashiwabara et al. Kashiwabara:2014 use association rule mining to identify verbs that might be good candidates for use in method names. Abebe Abebe:2013 uses an ontology that models the word relationships within a piece of software. Saeed et al. Saeed:2023 vectorize methods based on metrics and use the K-Nearst Neighbors algorithm with these vectors, and a large data set of methods, to recommend method names. Allamanis et al. Allamanis:2015 introduce a novel language model called the Subtoken Context Model. There has also been work to reverse engineer data types from identifiers malik2019 ; Hellendoorn:2018 . One thing these approaches have in common is the use of frequent tokens and source code context to try and generate high-quality identifier names (or understand their behavior for the purpose of generating types). There is a lot of work in this subfield, but the contrast to our work remains the same for all of them: These approaches aim to predict strong identifier names based on history. Our approach can help, since an understanding of common naming structures can support filtering out names that are inappropriate based on their grammatical structure; teach AI-based approaches how to optimize the identifiers they generate, or at least avoid using bad grammar structure; or help reverse-engineer the semantics of an identifier name based on its grammatical properties. In addition, automated name generation approaches cannot teach us very much about naming practices on their own, or help us formalize our understanding of strong naming structures, and how those can be taught in a classroom. Thus, our work is novel, and complementary to identifier name generation approaches.
7.5 Software Ontology Creation Using Identifier Names
A lot of work has been done in the area of modeling domain knowledge and word relationships by leveraging identifiers AbebeDomainConcepts ; RatiuProgramsKnowledgeBases ; RatiuConceptMapping ; Deissenboeck06aunified ; falleriWordnet . Abebe and Tonella AbebeDomainConcepts analyze the effectiveness of information retrieval-based techniques for filtering domain concepts and relations from implementation details. They show that fully automated techniques based on keywords or topics have low performance but that a semi-automated approach can significantly improve results. Falleri et al., present a way to automatically construct a wordnet-like miller1995wordnet identifier network from software. Their model is based on synonymy, hypernymy and hyponymy, which are types of relationships between words. Synonyms are words with similar or equivalent meaning; hyper/hyponyms are words which, relative to one another, have a broader or more narrow domain (e.g., dog is a hyponym of animal, animal is a hypernym of dog). Ratiu and Deissenboeck RatiuConceptMapping present a framework for mapping real world concepts to program elements bi-directionally. They use a set of object-oriented properties (e.g., isA, hasA) to map relationships between program elements and string matching to map these elements to external concepts. This extends two prior works of theirs: one paper on a previous version of their metamodel Deissenboeck06aunified and a second paper on linking programs to ontologies RatiuProgramsKnowledgeBases . Many of these approaches need to split and analyze words found in an identifier in order to connect these identifiers to a model of program semantics (e.g., class hierarchies). All of these approaches rely on identifiers.
Many software word ontologies use meta-data about words to understand the relationship between different words. There is a synergistic relationship between the work we present here and software ontologies, since stronger ontologies can help us generate and study grammar patterns effectively, and the CCID can help construct stronger software word ontologies. In particular, studying closed-category terms helps strengthen the metadata used to generate an ontology that seeks to map how words are related to one another, or code behavior.
7.6 Identifier Structure and Semantics Analysis
Liblit et al. Liblit06cognitiveperspectives discuss naming in several programming languages and makes observations about how natural language influences the use of words in these languages. Schankin et al. Schankin:2018 focus on investigating the impact of more informative identifiers on code comprehension. Their findings show the advantage of descriptive, compound identifiers over short single-word ones. Hofmeister et al Hofmeister:2017 compared comprehension of identifiers containing words against identifiers containing letters and/or abbreviations. Their results show that when identifiers contained only words instead of abbreviations or letters, developer comprehension speed increased by 19% on average. Lawrie et al. Binkley2006 did a study and used three different “levels” of identifiers. The results show that full-word identifiers lead to the best comprehension compared to the other levels studied. Butler’s work  butler2010exploring extends their previous work on Java class identifiers Butler:2009 to show that flawed method identifiers are also associated with low-quality code according to static analysis-based metrics. These papers primarily study the words found in identifiers and how they relate to code behavior or comprehension rather than word metadata (e.g., PoS).
Caprile and Tonella tonella:1999 analyze the syntax and semantics of function identifiers. They create classes which can be used to understand the behavior of a function; grouping function identifiers by leveraging the words within them to understand some of the semantics of those identifiers. While they do not identify particular grammar patterns, this study does identify grammatical elements in function identifiers, such as noun and verb, and discusses different roles that they play in expressing behavior both independently, and in conjunction, using the classes they propose. They also used the classes identified in this previous work to propose methods for restructuring program identifiers tonella:2000 . Fry and Shepherd Shepherd:2007 ; fry:2008 study verb-direct objects to link verbs to the natural-language-representation of the entity they act upon, in order to assist in locating action-oriented concerns. The primary concern in this work is identifying the entity (e.g., an object) which a verb is targeting (e.g., the action part of a method name).
Høst and Østvold study method names as part of a line of work discussed in Høst’s dissertation hostdissertation . This line of work starts by analyzing a corpus of Java method implementations to establish the meanings of verbs in method names based on method behavior, which they measure using a set of attributes which they define HostLexicon . They automatically create a lexicon of verbs that are commonly used by developers and a way to compare verbs in this lexicon by analyzing their program semantics. They build on this work in hostphrasebook by using full method names, which they refer to as phrases, and augment their semantic model by considering a richer set of attributes. The outcome is that they were able to aggregate methods by their phrases and come up with the semantics behind those phrases using their semantic model, therefore modeling the relationship between method names and method behavior. The phrases they discuss are similar to the general grammar patterns studied in our prior work Newman_GP . They extend this use of phrases by presenting an approach to debug method names Host:2009 . In this work, they designed automated naming rules using method signature elements. They use the phrase refinement from their prior paper, which takes a sequence of PoS tags (i.e., phrases) and concretizes them by substituting real words. (e.g., the phrase $<$verb$>$-$<$adjective$>$ might refine to is-empty). They connect these patterns to different method behaviors and use this to determine when a method’s name and implementation do not match. They consider this a naming bug. Finally, in Hst2011CanonicalMN , Høst and Østvold analyzed how ambiguous verbs in method names makes comprehension of Java programs more difficult. They proposed a way to detect when two or more verbs are synonymous and being used to describe the same behavior in a program; hoping to eliminate these redundancies as well as increase naming consistency and correctness. They perform this detection using two metrics which they introduce, called nominal and semantic entropy. Høst and Østvold’s work focuses heavily on method naming patterns; connecting these to the implementation of the method to both understand and critique method naming.
Butler butler:2011 studied class identifier names and lexical inheritance, analyzing the effect that interfaces or inheritance has on the name of a given class. For example, a class may inherit from a super class or implement a particular interface. Sometimes this class will incorporate words from the interface name or inherited class in its name. His study builds on work by Singer and Kirkham singer:2008 , who identified a grammar pattern for class names of (adjective)* (noun)+ and studies how class names correlate with micro patterns. Among Butler’s findings, he identifies a number of grammar patterns for class names: (noun)+, (adjective)+ (noun)+, (noun)+ (adjective)+ (noun)+, (verb) (noun)+ and extends these patterns to identify where inherited names and interface names appear in the pattern. The same author also studies Java field, argument, and variable naming structures butler:2015 . Among other results, they identify noun phrases as the most common pattern for field, argument, and variable names. Verb phrases are the second most common. Further, they discuss phrase structures for boolean variables; finding an increase in verb phrases compared to non-boolean variables. Olney Olney2016 compared taggers for accuracy on identifiers, but only on Java method names which were curated to remove ambiguous words (e.g., abbreviations).
Binkley et al Binkley:2011 studied grammar patterns for attribute names in classes. They come up with four rules for how to write attribute names: 1) Non-boolean field names should not contain a present tense verb, 2) field names should never only be a verb, 3) field names should never only be an adjective, and 4) boolean field names should contain a third-person form of the verb “to be” or the auxiliary verb “should”. Al Madi naser:2023 created a tool for performing lexical analysis of identifier names based on phonological, semantic, and orthographic similarity. Techniques that normalize identifiers, such as the one presented by Jingxuan Jingxuan:2023 , or by Hill hillamap can help make generating grammar patterns easier by expanding abbreviations into full words that a tagger can recognize more accurately. Aman et al. Aman:2024 studied confusing variable pairs, which are variables with very similar names, to understand how/if they are changed over time, and how pervasive they are.
None of the projects in this subsection deal specifically with closed-category grammar patterns, or even terms that fall within a closed PoS category. Many of them, particularly the work on PoS taggers, on grammar patterns in differing contexts, normalizing identifier names, and on grammatical anti-patterns, are likely mutually-synergistic to our work. This is because a stronger understanding of closed-category terms/patterns, and how they relate to program behavior, can help support the style of analysis these works leverage.
8 Discussion and Future Work
Our analysis focuses deeply on the semantic roles of closed-category terms, including determiners, prepositions, conjunctions, and digits; as well as the behavioral codes they imply. By applying axial and selective coding to a hand-curated dataset, we uncovered patterns of usage that go beyond surface grammar and reveal how developers embed control flow, intent, and logical relationships into naming conventions.
These findings reveal that closed-category terms serve as purposeful, behaviorally expressive units of meaning that often directly map to behavioral constructs in code. Developers use these compact grammatical forms as cognitive shortcuts to signal structural distinctions, state transitions, and operational logic. This strategic use of language suggests that these terms are important to understand and study.
  1. 1.
    Closed-category terms convey rich and distinct behavioral semantics. We examined the unique roles that each type of closed-category term tends to serve. For instance, determiners often reflect selection (e.g., someItem), temporality (e.g., nextNode), or negation (e.g., noCache); conjunctions signal alternatives, conjunctional guards, or composite actions (e.g., save_and_close, key_or_iv); and digits most commonly serve to distinguish, enumerate, or encode system-specific roles (e.g., arg1, Neo4j). These behavioral distinctions have been formalized in our axial coding schema, which can serve as a foundation for recommendation systems, naming audits, or comprehension studies.
  2. 2.
    Closed-category semantics adapt to programming context in ways distinct from natural language. Prepositions that typically express motion or containment in English (e.g., to, from, on) are disproportionately found in function names, where they signal behavior, transformation, or event triggers. Conversely, determiners like no, this, and next are rare in class names but common in parameters or attributes, where positional reasoning and state scoping are essential. This contextual distribution suggests that developers adapt the semantics of closed-category terms to the structural and functional roles of identifiers, shaping how behavior is encoded in code.
  3. 3.
    Closed-category terms serve as lexical scaffolds for program comprehension. These terms often compress multiple behavioral or structural ideas into a compact name. Whether flagging control flow (e.g., ifEnabled), encoding default values (noVal), or signaling transformation (toString), they act as micro-annotations; embedding traceability, logic, or domain-specific cues directly into identifiers. This use of grammar-as-guidance offers a potentially promising direction for tools aimed at improving name clarity and developer communication.
  4. 4.
    Grammar patterns help us understand form and function. Combining axial and selective codes with grammar patterns reveals both the entities involved (open-category terms) and the behavioral relationship between them (closed-category terms). When the closed term appears at the beginning of a pattern (e.g., DT NM N, P NM N), it typically forms a unary relation, modifying a single operand (e.g., temporal status or type reinterpretation). When it appears between open terms (e.g., N P N, N CJ N), it often encodes a binary relation—linking two operands through behavior or logic (e.g., data flow or disjunction). We also note that function identifiers with a final closed term (e.g., sendTo, readFrom) may also express a binary relation, where the second operand is supplied as a parameter rather than named directly. This is consistent with observations made during data analysis in our prior work Newman_GP .
These findings have practical implications for tool builders and educators interested in improving naming support. For example, grammar patterns could be integrated into static analysis or IDE plugins to provide optional, contextual suggestions; highlighting when an identifier follows an uncommon structure or when the pattern contrasts with its code context. This does not imply the name is incorrect, but it may prompt a developer or reviewer to reflect on whether the chosen pattern aligns with the intended semantics. For instance, encountering a pattern like N CJ N (e.g., dataOrLogger) in a constant declaration could trigger a soft prompt: “This naming pattern is rare in this context—consider whether it clearly communicates its role.”
Grammar patterns can also be used to scaffold naming suggestions in LLM-driven tools. Instead of generating identifiers purely from task descriptions, models could be prompted to produce names that instantiate common closed-category structures (e.g., DT NM N, P N), which are frequently associated with behavioral semantics. This approach may help align completions with human naming conventions, while still allowing flexibility in term choice.
Beyond tooling, grammar pattern awareness can enhance educational workflows. Instructors could use pattern frequency and semantics to illustrate naming “idioms,” helping students understand how experienced developers encode behavior through compact syntactic forms. Similarly, code review tools might use grammar pattern summaries to draw attention to unconventional naming constructs, offering reviewers an additional signal without enforcing rigid standards.
In future work, we plan to evaluate these ideas empirically: measuring whether adherence to common patterns improves comprehension, how grammar scaffolding affects naming quality in LLM-generated identifiers, and whether tools that surface grammar patterns can meaningfully assist developers. In addition, it would be interesting to perform studies simialar to Schankin Schankin:2018 and Hofmeister Hofmeister:2017 ’s work on how the descriptiveness of names influence comprehension; or Arnaoudova Arnaoudova:2013 , and HostHost:2009 who looked at how code behavior and naming structure could be used to measure name quality. Specifically, we could study how closed-category terms change or augment the outcomes of their studies.
9 Threats to Validity
Construct Validity: This study is conducted on a manually annotated dataset of 1,275 identifiers containing closed-category terms, the largest of its kind at the time of writing. A potential threat lies in the completeness of our closed-category term list: we relied on a predefined lexicon (Section 4), meaning novel or unlisted terms may be absent from the dataset. However, their absence would likely expand our results rather than refute them. Our identifier sample is restricted to production code, and although we excluded known test files, developers may occasionally include test logic in production files. To mitigate this, we manually reviewed each identifier and its source context. Furthermore, while we used file extensions to distinguish C (.c, .h) from C++ (.cpp, .hpp), these conventions are not absolute. We addressed this by manually validating the source language of each identifier.
Internal Validity: Abbreviations within identifiers were not expanded, which may have caused occasional misinterpretation by annotators. However, annotators had access to the surrounding source code, reducing the risk of misannotation. Grammar pattern tagging and axial coding for each closed-category term were both subject to cross-annotation by three independent annotators and evaluated using Fleiss’ Kappa to assess agreement. We used a grounded-theory approach to develop our behavioral codes. Four coders participated in open and axial coding; during the selective coding phase, one coder proposed all selective codes, which were then validated and refined collaboratively by the other three through discussion until thematic saturation was reached.
We used statistical methods to examine correlations between closed-category terms and contextual variables. We performed two chi-square tests: one to assess correlation between closed-category part-of-speech categories (e.g., Determiner, Preposition) and programming language (Java, C, C++), and another for their correlation with code context (Attribute, Function, Declaration, Parameter, Class), derived automatically via srcML collard:2016 . We applied Bonferroni correction to account for multiple comparisons. A threat to internal validity is the assumption of independence in the chi-squared test. If violated, some significance values may be distorted. However, the primary insights of RQ1, which focus on behavioral coding through qualitative analysis, are unaffected by this statistical assumption.
External Validity: Our data includes identifiers from C, C++, and Java, three widely used languages with similar syntactic and object-oriented paradigms. While this helps reduce language-specific bias, our findings may not generalize to other paradigms such as functional or logic-based languages, where naming conventions and code contexts may differ significantly.
Mitigation Strategies: To ensure transparency and reproducibility, the dataset will be made publicly available (Section 11). Annotators were allowed to inspect source code when labeling identifiers, and each identifier was independently annotated twice. Grammar patterns and axial codes were validated by multiple annotators, with inter-rater agreement assessed using Fleiss’ Kappa. We selected a representative sample from 30 software systems, sized to meet a 95% confidence level with a 5% confidence interval. Code context was derived automatically using srcML. Finally, to evaluate whether closed-category term usage varies by domain, we curated a domain-specific dataset (e.g., compilers, databases, networking tools) and compared it against a general-purpose set selected without regard to domain. We applied a Mann-Whitney U test to compare term frequencies between these groups, normalizing by lines of code to control for system size.
10 Conclusions
This paper presents a detailed empirical study of closed-category terms in identifier names, highlighting how they function semantically across a wide range of software artifacts. Our contributions include:
  1. 1.
    The CCID dataset: A new, part-of-speech–annotated dataset of identifier names containing closed-category terms, released with this paper. It supplements prior datasets focused on open-class lexical items, enabling more nuanced research into naming semantics.
  2. 2.
    A dual-level coding framework: We introduce a combined selective and axial coding scheme to interpret the semantics of determiners, prepositions, conjunctions, and digits in identifiers. This framework maps grammar structure to conceptual behavior in a way not previously formalized, and provides a solid basis for future research and development of naming practices.
  3. 3.
    Insights into the distinct semantic roles of closed-category words: Our study shows how programming diverges from natural language in its use of terms like next, both, and or, and how these terms reflect functional intent embedded in code structure.
These findings have practical implications for code review, refactoring, and intelligent naming support. Automated tools can leverage our work to suggest context-appropriate names, detect inconsistent patterns, or provide just-in-time feedback during development. The insights we have uncovered can be packaged and curated by educators to teach semantic clarity in naming, moving beyond generic best practices to behavior-specific guidance.
We also outline several promising directions for future research:
  1. 1.
    Further validation of the behavioral categories we propose, especially through comprehension studies that measure how different naming structures affect developer understanding.
  2. 2.
    Experimental work to test which closed-category naming patterns improve or hinder program comprehension, while controlling for code context and developer experience.
  3. 3.
    Integration of our findings into intelligent development tools, including naming recommendation systems, refactoring support, and automated code reviewers capable of detecting semantic mismatches or naming anti-patterns.
This work demonstrates that closed-category terms are not linguistic noise; they are deliberate, behaviorally meaningful tools in the software naming arsenal. By mapping their roles across grammar patterns and program contexts, we provide a foundation for future studies in naming semantics and for tools that support naming literacy. We believe our findings raise the question: Should closed-category terms be used more often in naming? It is clear that they are used purposefully, and we believe that their use does help comprehension, but in which situations is that true? And how can we detect when such terms should be used, as opposed to another naming pattern? We believe these findings can help guide tool builders and researchers in novel, fruitful directions that formalize and improve naming practices among developers.
11 Data Availability
We have created a repository that contains the data and scripts needed to generate the numbers and statistical analysis from the RQs. The scripts are in the scripts directory and need only to be run (they take no arguments). The data directory contains all of the annotation data, both broken down by closed category; each file is named after the category it contains, and amalgamated in a single file (called Tagger Open Coding). This repository can be found at this link [^5] https://github.com/SCANL/closed_category_emse_analysis_scripts.
12 Competing Interests
We have no competing/conflicting interests to report